NOVEL PATTERNS OF TORSION - INVERSION TUNNELING AND TORSION - ROTATION COUPLING IN THE ν_{11} CH - STRETCH REGION OF CH_{3}NH_{2}

\underline{MAHESH B DAWADI}, SYLVESTRE TWAGIRAYEZU, C. MICHAEL LINDSAY,^{a} AND DAVID S. PERRY, Department of Chemistry, The University of Akron, OH 44325-3601; LI-HONG XU, Department of Physics, Centre for Laser, Atomic and Molecular Studies (CLAMS) University of New Brunswick, Saint John, New Brunswick, Canada E2L 4L5.

The high-resolution infrared spectrum of CH_{3}NH_{2} has been recorded using slit-jet absorption spectroscopy in the ν_{11} asymmetric CH-stretch region (2965 to 3005 cm^{−1}) with resolution of 0.003 cm^{−1}. The 580 lines, assigned by ground state combination differences, represent 27 subbands with |K′| ≤ 2 for the A, B, E_{1} and E_{2} symmetries. Several of the observed subbands are split by perturbations. The analysis of spectrum shows that the patterns of the torsion-inversion tunneling splittings are qualitatively different from the ground state. In addition, the low - J splittings between |K′| = +|K′| and −|K′| are greatly reduced relative to the ground state in both the E_{1} and the E_{2} species, indicating that torsion-rotation coupling is suppressed in the ν_{11} CH-stretch excited state.

aPresent address: U.S. Air Force Research Laboratory, 2306 Perimeter Rd, Eglin AFB, FL 32542-5910